OpenStudy (rianiscool):

need help asap will give a meal Solve using the guess-and-test strategy. The math club has 57 members who are in sixth or seventh grades. There are twice as many sixth graders as seventh graders. How many sixth graders are in the club? A. 38 B. 37 C. 19 D. 18

OpenStudy (anonymous):

57 = 3y 19 = y There are y (19) seventh graders and 2y (38) sixth graders. For guess-and-check, you must simply add the number times the number divided by two and see if the answer is the total count of members (57). For instance, for (A): ([38] + [38/2]) = (38 + 19) = 57, so (A) is correct.

OpenStudy (rianiscool):

thank you